OpenStudy (anonymous):

Solve the system by substitution -4.5x-2y=-12.5 3.25x-y=-0.75

OpenStudy (anonymous):

Do you know how to isolate a variable?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

To get the y in the bottom equation isolated, subtract the 3.25x to the other side. This will leave you with -y=.75-3.25x. Then multiply everything by -1 to get the y positive. This will leave you with y=-.75+3.25x. After that put the y=-.75+3.25x into the top equation which will get you -4.5x-2(-.75+3.25x)=-12.5 Should be easy from there

OpenStudy (anonymous):

I'm positive I got this wrong, but is it -1.25x-1.25=-12.5?

OpenStudy (phi):

*** Solve the system by substitution -4.5x-2y=-12.5 3.25x-y=-0.75 *** "by substitution" means use one of the equations, and "solve" for one of the variables in this case, the 2nd equation is easiest to use..if we solve for "y" 3.25x-y=-0.75 first, we can add +0.75 to both sides, like this 3.25x+0.75 -y=-0.75+0.75

OpenStudy (phi):

3.25x+0.75 -y=-0.75+0.75 the right side simplifies to 0. we have 3.25x+0.75 -y= 0 now add +y to both sides: 3.25x+0.75 -y + y=0+y that simplifies to 3.25x+0.75 = y or y = 3.25x+0.75

OpenStudy (phi):

now you have what y is equal to: y = 3.25x+0.75 use that expression in the first equation -4.5x-2y=-12.5 <-- replace y with (3.25x+0.75) you get -4.5x-2(3.25x+0.75)=-12.5 now to solve for x, you should think of this as -4.5x + -2(3.25x+0.75)=-12.5 The first step is "distribute" the -2. that means multiply -2 times each term inside the parens. Can you do that ?

OpenStudy (anonymous):

@Phi do I subtract -4.5x from -6.5x? Or rather add them?

OpenStudy (phi):

First, after you distribute the -2 , you should get -4.5x -6.5x -1.5 = -12.5 to avoid confusion, you should change subtraction to "adding negative numbers (or letters)" in other words, write this as -4.5x + -6.5x + -1.5 = -12.5 which means the same thing, but makes it more clear (to me anyways) that you add the two negative terms to get an even more negative term... can you simplify ?

OpenStudy (anonymous):

-4.5x+-6.5x+-1.5=-12.5

OpenStudy (phi):

you should combine the "like terms" It is easier to see with "nice" numbers: if you had 1x + 2x that means 1 x plus 2 more x's which is 3 x's or 3x or if you had -1x - 2x that means -1 x take away 2 more x's gives -3 x's or -3x in your case you have -4.5x+-6.5x can you combine those two terms ?

OpenStudy (anonymous):

I had actually combined them in my head, I just don't know why I didn't type it. -11x+-1.5=-12.5

OpenStudy (phi):

too bad, I could have saved all that typing. now add +1.5 to both sides

OpenStudy (anonymous):

-11x+-1.5+1.5=-12.5+1.5 -11=-11

OpenStudy (phi):

I hope you mean -11x = -11

OpenStudy (phi):

now divide both sides by -11

OpenStudy (anonymous):

1

OpenStudy (anonymous):

1=1

OpenStudy (phi):

you should write out the whole equation (a good habit, which is helpful to not getting lost for complicated problems) -11x = -11 \[ \frac{-11}{-11} x = \frac{-11}{-11} \\ x=1 \] now you need to find y from way up above, we have this equation y = 3.25x+0.75 can you find y when x = 1 ?

OpenStudy (anonymous):

3.25(1)+0.75=4 y=4

OpenStudy (phi):

yes, that looks good. you can always check your answer by replacing x with 1 and y with 4 in the original equations and make sure they are true.

OpenStudy (anonymous):

Thank you!!